# FT2TF: First-Person Statement Text-To-Talking Face Generation

### v1

@@ Abstract
Talking face generation has gained immense popularity in the computer vision community, with various applications including AR/VR, teleconferencing, digital assistants, and avatars. Traditional methods are mainly audio-driven ones which have to deal with the inevitable resource-intensive nature of audio storage and processing. To address such a challenge, we propose FT2TF – First-Person Statement Text-To-Talking Face Generation, a novel one-stage end-to-end pipeline for talking face generation driven by first-person statement text. Moreover, FT2TF implements accurate manipulation of the facial expressions by altering the corresponding input text. Different from previous work, our model only leverages visual and textual information without any other sources (e.g. audio/landmark/pose) during inference. Extensive experiments are conducted on LRS2 and LRS3 datasets, and results on multi-dimensional evaluation metrics are reported. Both quantitative and qualitative results showcase that FT2TF outperforms existing relevant methods and reaches the state-of-the-art. This achievement highlights our model’s capability to bridge first-person statements and dynamic face generation, providing insightful guidance for future work.
Under Review
@@ 1 Introduction
Talking face generation, with the goal of creating a sequence of faces to simulate the appearance of a person talking, has surged in significance amid the metaverse boom and technological advancements. This pursuit has gained substantial popularity and importance within the computer vision community, influencing applications including augmented reality [16], virtual reality [9], computer games [30, 29], teleconferencing [7, 3], video dubbing [11, 18], movie special effects [43], and the development of digital assistants [33] and avatars [9, 49].
Traditional methods for talking face generation mainly rely on audio input, which leverages visual and audio cues. These methods, exemplified by works like [28, 49, 48, 17, 46], show impressive performance. However, they come with a significant drawback: the need for high-quality audio, leading to strict requirements on storage space, data transmission, and recording environments, as highlighted in [34]. Addressing such challenges holds significant value not only for the scientific community but also for the industry, aiming to mitigate the environmental impact of AI technologies  [32]. Consequently, one crucial question has arisen: Is it feasible to substitute audio with text inputs while ensuring detailed facial expressions of generated frames?
@@
There exist two primary types of text-driven face generation methods. The first involves using third-person text descriptions where a facial expression is described from an external perspective. These descriptions are leveraged to generate realistic faces using end-to-end GAN architectures [13, 25, 41]. The second path employs first-person statements which are the speaker’s spoken captions captured in text form. While this approach aligns more closely with talking face generation, it has not been thoroughly explored. Previous methods mainly conduct a two-stage framework including text-to-speech and speech-to-face procedures [27, 45]. These approaches generate audio waveforms from text statements and further generate faces based on synthesized audio [31]. Therefore, methods that follow a one-stage end-to-end pattern to generate talking faces driven by first-person statements remain unestablished. Based on this, our objective is to revolutionize the synthesis of coherent and natural talking face videos through first-person statements, encompassing well-synchronized lip movements, accurate cheek actions, detailed textures, and overall facial expressions, as illustrated in Figure 1.
To overcome the challenges above, we propose First-Person Statement Text-To-Talking Face Generation (FT2TF) – a one-stage end-to-end pipeline that effectively generates high-quality talking faces from visual-textual cross-modality. Different from previous work, our pipeline encapsulates the entire face generation process without any other sources (e.g. audio, landmarks, poses) being used during inference.
@@
  1. 1.
    End-to-end Continuous-Frame Talking Face Generation. We propose FT2TF – an effective one-stage pipeline that generates realistic talking faces by integrating visual and textual input. FT2TF evidently enhances face naturalness, synchronizes lip movements, and improves detailed facial expressions.
  2. 2.
    Effective Facial Expression Manipulation. FT2TF enables manipulations of the speaker’s facial expressions and lip movements by altering corresponding input text.
  3. 3.
@@ 2.1 Audio-Driven Talking Face Generation
Whole Face Generation. Zhou et al. [49] disentangles audio sources into content and speaker embeddings, followed by an LSTM-based module for landmark prediction to generate the talking head. Based on that, models that use input sources other than audio have been proposed. PC-AVS [48] utilizes the pose source to control the head pose of the synthesized faces, while EAMM [17] uses pose sequence, emotion source, and audio ground truth to generate the target video manipulated by different emotion guidances. In contrast to existing methods that involve multiple input resources for face manipulation, our model only uses text and reference face frames as the input source. This ensures the efficient generation of high-quality facial representations, free from input requirements such as constraints of audio quality.
@@ 2.2 Text-Driven Talking Face Generation
The exploration of talking face generation based on first-person statements has not been well-established. Initially, models using Variational Autoencoder (VAE) [20, 24, 37, 40] and CRNN [4, 35] are designed to generate video based on text input. However, the generated frames by these models are unrealistic and over-smoothing, especially in long-sequence video generation tasks [40, 26].Text-driven talking face generation has mainly been divided into two directions: using first-person statements or third-person descriptions. Methods driven by third-person descriptions [13, 25, 41] aim to learn the mapping between objective text descriptions and synthesized video. Most relevant approaches using first-person statements are constructed in a two-stage architecture with two separate steps: converting text to speech and then generating faces from the speech [21, 22, 42, 27, 45]. Consequently, one-stage methods that generate talking faces directly from first-person text captions remain unexplored. Addressing this research gap, our model is constructed as a one-stage end-to-end framework that directly leverages reference talking face frames and first-person text caption to generate high-quality talking faces efficiently. Since relevant exploration has not been widely established, our solution aims to bridge first-person text and dynamic face synthesis to provide insightful guidance for future work.
3 Method
@@
We address the challenging problem of generating realistic talking face videos from first-person text descriptions. Formally, given $M$ text caption tokens $X^{t}_{1:M}$ as the complete sentence and the initial $k$ talking face frames $X^{v}_{1:k}$ as the input, our objective is to synthesize the subsequent talking face frames $X^{v}_{k+1:N}$. In this context, $t$ represents the textual content, and $v$ represents the visual content. The text captions $X^{t}_{1:M}$ and the talking face frames $X^{v}_{1:N}$ are aligned to correspond with each other.
@@
To tackle the problem, we present a novel pipeline – FT2TF: First-Person Statement Text-to-Talking Face Generation, as illustrated in Figure 2. We use two specialized text encoders – the Global Emotion Text Encoder captures overall emotional tone from complete text caption $X^{t}_{1:M}$, and the Linguistic Text Encoder encodes semantic compression from the reference text caption $X^{t}_{m+1:M}$ which corresponds to frames $X^{v}_{k+1:N}$. Alongside these, a Visual Encoder is employed to extract visual features from reference frames $X^{v}_{1:k}$. Inspired by [23], we design and utilize a Multi-Scale Cross-Attention Module that integrates Global and Local Cross-Attention modules to align features from two modalities into the same distribution, aiming to ensure a comprehensive visual-textual fusion across varying scales and feature dimensions. The resulting visual-textual representations are fed to a Visual Decoder to synthesize talking face frames $X^{v}_{k+1:N}$. Implementation details and structural insights are provided in the supplementary material.
@@ 3.1 Multimodal Encoders
Global Emotion Text Encoder. The Global Emotion Text Encoder harnesses the capabilities of the pre-trained Emoberta [19] model. Emoberta serves as the cornerstone for encoding the overarching emotional context embedded within textual descriptions. In light of this, it presents itself as a prime contender for global attention. By utilizing the entire face frames, we can direct our exclusive focus toward text emotions and their face representations.
Linguistic Text Encoder. We adopt to use the pre-trained GPT-Neo [10] as the Linguistic Text Encoder. In our choice to employ this model, we find motivation in the distinctive characteristics of the GPT-Neo, as opposed to the Emoberta [19]. GPT-Neo has undergone training using a vast language dataset, which empowers it with enhanced linguistic compression for every statement conveyed by the talking face. Consequently, we intend to narrow our attention solely to the region of the lips, aligning with the phonetic articulation of each word, as can be seen in Figure 2(c).
@@
3.2 Multi-Scale Cross-Attention and Visual Decoder
Given the encoded textual ($F^{t}_{emo}$, $F^{t}_{ling}$) and visual features ($F^{v}$), we design a Multi-Scale Cross-Attention module that integrates Global and Local Cross-Attention modules. Specifically, features extracted from the Visual Encoder, $F^{v}$, are used in conjunction with the textual features from the Global Emotion Text Encoder, $F^{t}_{emo}$, as the input to the Global Cross-Attention. In this context, $F^{v}$ serves as both the key ($K$) and value ($V$), while $F^{t}_{emo}$ functions as the query ($Q$). A similar process occurs with the Local Cross-Attention where $F^{v}$ serves as $K$ and $V$ and the output of Linguistic Text Encoder, $F^{t}_{ling}$, acts as $Q$.
@@ 3.2 Multi-Scale Cross-Attention and Visual Decoder
Multi-Scale Cross-Attention. The collective use of Global and Local Cross-Attention ensures a balanced fusion of global emotional context and local linguistic features. As corroborated by Section 4, this multimodal fusion mechanism empowers FT2TF to generate talking face videos with natural lip synchronization and emotionally expressive facial details, resulting in more realistic and engaging outputs.
@@ 3.3 Pipeline Training
Three key loss functions are constructed for the pipeline training.
Generation Loss. The generation loss, $L_{gen}$, measures the pixel-wise discrepancy between the generated frames and the ground truth ones. This loss primarily focuses on supervising the accurate generation of pixel-level details to ensure the fidelity of the synthesized frames, as shown below:
$L_{gen}=\frac{1}{N}\sum_{i=1}^{N}\left\|X^{v}_{k+i}-O^{tv}_{k+i}\right\|_{1}$ (1)
@@
$L_{syn}=\frac{1}{N}\sum_{i=1}^{N}SyncNet(O^{tv}_{k+i},A_{k+i})$ (2)
Discriminator Loss. The discriminator loss [12], $L_{disc}$, aims to guide the encoder-decoder process to produce more realistic and coherent images. It operates by utilizing a discriminator that provides binary predictions $O_{pred}$ based on whether the input is the generated one or the ground truth. Specifically, the corresponding labels $Y_{disc}$ are assigned the label 1 for generated data and 0 for ground truth, yielding a binary cross-entropy loss:
$\begin{split}L_{disc}=-\frac{1}{N}\sum_{i=1}^{N}(&Y_{disc}\cdot\log(O_{pred})\\ &+(1-Y_{disc})\cdot\log(1-O_{pred}))\end{split}$ (3)
@@
$L_{total}=\lambda_{1}L_{gen}+\lambda_{2}L_{syn}+\lambda_{3}L_{disc}$ (4)
@@ 4 Experiment
Method Dataset Generation Input PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$ LipLMDnormal-$\downarrow$ CSIMnormal-$\uparrow$
@@
4.1 Datasets
  • LRS2 [1]: The dataset consists of 48,164 video clips collected from BBC outdoor broadcasts, and is split into the train, validation, and test sets containing 45,839, 1,082, and 1,243 video clips, respectively. Each clip is paired with a text caption containing up to 100 characters.
  • LRS3 [2]: The LRS3 dataset is constructed from over 400 hours of 5,594 indoor TED/TEDx shows with 151,819 videos, along with the corresponding text captions. Specifically, it consists of 32,000 train/val utterances and 1,452 test ones.
While the two datasets share a similar name, it is essential to clarify that there is no overlap between them. Videos in these two datasets differ significantly in terms of the scenes captured, collection methods, character actions, environmental conditions, and other factors. Supplementary material provides data pre-processing details.
4.2 Comparison Methods
We conduct a comprehensive evaluation of our method against state-of-the-art approaches for talking face generation. Due to limited research in first-person text-driven talking face generation, we demonstrate the effectiveness and efficiency of our model by comparing it with audio-driven models on LRS2 and LRS3 datasets. To provide a clear and comprehensive assessment, we categorize these models based on two criteria:
  • Generation Task: All methods are categorized into either Lip Region or Whole Face, with the former generating only the lip region and the latter generating the entire face.
  • Modality: All methods are classified based on the input modality, including Audio+Video (A+V), Text+Video (T+V), Audio+Video+Pose/Landmarks (A+V+P/L).
4.3 Evaluation Metrics
We employ multiple evaluation metrics across various criteria, ranging from evaluating video synthesis quality to evaluating facial expression representation. To evaluate the similarity between generated results and ground truth ones, we apply Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) [38]. In addition, we utilize Learned Perceptual Image Patch Similarity (LPIPS) [44] and Frechet Inception Distance (FID) [14] for assessing feature-level similarity. Considering the identity preservation, we leverage cosine similarity (CSIM) between identity vectors extracted by the ArcFace [8] network. Moreover, aligning with the approach outlined in [39], we measure the normalized lip landmarks distance (LipLMD) between generated and ground truth frames to evaluate lip synchronization quality.

### v2

@@ Abstract
Talking face generation has gained immense popularity in the computer vision community, with various applications including AR, VR, teleconferencing, digital assistants, and avatars. Traditional methods are mainly audio-driven, which have to deal with the inevitable resource-intensive nature of audio storage and processing. To address such a challenge, we propose FT2TFFirst-Person Statement Text-To-Talking Face Generation, a novel one-stage end-to-end pipeline for talking face generation driven by first-person statement text. Different from previous work, our model only leverages visual and textual information without any other sources (e.g., audio/landmark/pose) during inference. Extensive experiments are conducted on LRS2 and LRS3 datasets, and results on multi-dimensional evaluation metrics are reported. Both quantitative and qualitative results showcase that FT2TF outperforms existing relevant methods and reaches the state-of-the-art. This achievement highlights our model’s capability to bridge first-person statements and dynamic face generation, providing insightful guidance for future work.
2025 IEEE/CVF Winter Conference on Applications of Computer Vision (WACV)
@@ 1 Introduction
Talking face generation, with the goal of creating a sequence of faces to simulate the appearance of a person talking, has surged in significance amid the metaverse boom and technological advancements. This pursuit has gained substantial popularity and importance within the computer vision community, influencing applications including augmented reality [17], virtual reality [9], computer games [33, 32], teleconferencing [7, 3], video dubbing [11, 19], special effects for movies [46], and the development of digital assistants [36] and avatars [9, 52].
@@
Traditional methods for talking face generation mainly rely on audio input, which leverages visual and audio cues. These methods, exemplified by works like [31, 52, 50, 18, 49], show impressive performance. However, they come with a significant drawback: the need for high-quality audio. Compared to audio-driven methods, text-driven approaches offer unparalleled advantages, notably in terms of:
  • Data Storage: Text data storage requires far less space than audio data. For instance, storing the text of a 300-page article requires a mere 128 KB, while storing audio for the same content demands several gigabytes.
  • Data Transmission: Current audio-driven talking face generation methods necessitate high-quality audio, imposing stringent requirements on data transmission conditions and recording environments [37]. In contrast, text-driven talking face generation methods operate effectively with basic and reliable network transmission conditions.
Addressing such challenges holds significant value not only for the scientific community but also for the industry, aiming to mitigate the environmental impact of AI technologies  [35]. Consequently, one crucial question has arisen: Is it feasible to substitute audio with text inputs while ensuring detailed facial expressions of generated frames?
There exist two primary types of text-driven face generation methods. The first involves using third-person text descriptions where a facial expression is described from an external perspective. These descriptions are leveraged to generate realistic faces using end-to-end GAN architectures [13, 28, 44]. The second path employs first-person statements which are the speaker’s spoken captions captured in text form. While this approach aligns more closely with talking face generation, it has not been thoroughly explored. Previous methods mainly conduct a two-stage framework including text-to-speech and speech-to-face procedures [30, 48]. These approaches generate audio waveforms from text statements and further generate faces based on synthesized audio [34]. Therefore, methods that follow a one-stage end-to-end pattern to generate talking faces driven by first-person statements remain unestablished. Based on this, our objective is to revolutionize the synthesis of coherent and natural talking face videos through first-person statements, encompassing well-synchronized lip movements, accurate cheek actions, detailed textures, and overall facial expressions, as illustrated in Figure 1. To overcome the challenges above, we propose First-Person Statement Text-To-Talking Face Generation (FT2TF) – a one-stage end-to-end pipeline that effectively generates high-quality talking faces from visual-textual cross-modality. Different from previous work, our pipeline encapsulates the entire face generation process without any other sources (e.g., audio, landmarks, poses) being used during inference.
@@
  • One-stage End-to-end Continuous-Frame Talking Face Generation. We propose FT2TF, a one-stage end-to-end pipeline that generates realistic continuous-frame talking faces by integrating visual and textual input.
  • Effective and Efficient Text-Driven Talking Face Generation. FT2TF utilizes only visual and textual information during inference. It enhances face naturalness, synchronizes lip movements, and improves detailed facial expressions while using fewer trainable parameters.
@@ 2.1 Audio-Driven Talking Face Generation
Whole Face Generation. Zhou et al. [52] disentangles audio sources into content and speaker embeddings, followed by an LSTM-based module for landmark prediction to generate the talking head. Based on that, models that use input sources other than audio have been proposed. PC-AVS [50] utilizes the pose source to control the head pose of the synthesized faces, while EAMM [18] uses pose sequence, emotion source, and audio ground truth to generate the target video manipulated by different emotion guidances. In contrast to existing methods that involve multiple input resources for face generation, our model only uses text and reference face frames as the input source. This ensures the efficient generation of high-quality facial representations, free from input requirements such as constraints of audio quality.
@@ 2.2 Text-Driven Talking Face Generation
The exploration of talking face generation based on first-person statements has not been well-established. Initially, models using Variational Autoencoder (VAE) [21, 26, 40, 43] and CRNN [4, 38] are designed to generate video based on text input. However, the generated frames by these models are unrealistic and over-smoothing, especially in long-sequence video generation tasks [43, 29]. Text-driven talking face generation has been divided mainly into two directions: using first-person statements or third-person descriptions. Methods driven by third-person descriptions [13, 28, 44] aim to learn the mapping between objective text descriptions and synthesized video. The most relevant approaches using first-person statements are constructed in a two-stage architecture with two separate steps: converting text to speech and then generating faces from the speech [22, 23, 45, 30, 48]. The latest work, TTFS [16], integrates text-to-speech and face-generation modules to synthesize talking face and audio from input text jointly. Consequently, one-stage methods that generate talking faces directly from first-person text captions remain unexplored [25, 27]. Addressing this research gap, our model is constructed as a one-stage end-to-end framework that directly leverages reference talking face frames and first-person text caption to generate high-quality talking faces efficiently. Since relevant exploration has not been widely established, our solution aims to bridge first-person text and dynamic face synthesis to provide insightful guidance for future work.
3 FT2TF Pipeline
@@
We address the challenging problem of generating realistic talking face videos from first-person text descriptions. Formally, given $M$ text caption tokens $X^{t}_{1:M}$ as the complete sentence and the initial $k$ talking face frames $X^{v}_{1:k}$ as input, our objective is to synthesize the subsequent talking face frames $X^{v}_{k+1:N}$. In this context, $t$ represents the textual content, and $v$ represents the visual content. The text captions $X^{t}_{1:M}$ and the talking face frames $X^{v}_{1:N}$ are aligned to correspond to each other.
@@
To tackle the problem, we present a novel pipeline – FT2TF: First-Person Statement Text-to-Talking Face Generation, as illustrated in Figure 2. We use two specialized text encoders – the Global Emotion Text Encoder captures overall emotional tone from the complete text caption $X^{t}_{1:M}$, and the Linguistic Text Encoder encodes semantic compression from the reference text caption $X^{t}_{m+1:M}$ which corresponds to frames $X^{v}_{k+1:N}$. Alongside these, a Visual Encoder is employed to extract visual features from reference frames $X^{v}_{1:k}$. Inspired by [24], we design and utilize a Multi-Scale Cross-Attention Module that integrates Global and Local Cross-Attention modules to align features from two modalities into the same distribution [39, 24], aiming to ensure a comprehensive visual-textual fusion across varying scales and feature dimensions. The resulting visual-textual representations are fed to a Visual Decoder to synthesize talking face frames $X^{v}_{k+1:N}$.
@@ 3.1 Multimodal Encoders
Global Emotion Text Encoder. The Global Emotion Text Encoder harnesses the capabilities of the pretrained Emoberta [20] model. Emoberta serves as the cornerstone for encoding the overarching emotional context embedded within textual descriptions. In light of this, it presents itself as a prime contender for global attention. Using the entire face frames, we can direct our exclusive focus toward text emotions and their face representations.
Linguistic Text Encoder. We adopt to use the pretrained GPT-Neo [10] as the Linguistic Text Encoder. In our choice to employ this model, we find motivation in the distinctive characteristics of the GPT-Neo, as opposed to the Emoberta [20]. GPT-Neo has undergone training using a vast language dataset, which empowers it with enhanced linguistic compression for every statement conveyed by the talking face. Consequently, we intend to narrow our attention solely to the region of the lips, aligned with the phonetic articulation of each word, as can be seen in Figure 3.
3.2 Multi-Scale Cross-Attention and Visual Decoder
Given the encoded textual ($F^{t}_{emo}$, $F^{t}_{ling}$) and visual features ($F^{v}$), we design a Multi-Scale Cross-Attention module that integrates Global and Local Cross-Attention modules. Specifically, features extracted from the Visual Encoder, $F^{v}$, are used in conjunction with the textual features from the Global Emotion Text Encoder, $F^{t}_{emo}$, as input to the Global Cross-Attention. In this context, $F^{v}$ serves as both the key ($K$) and the value ($V$), while $F^{t}_{emo}$ functions as the query ($Q$). A similar process occurs with Local Cross-Attention where $F^{v}$ serves as $K$ and $V$ and the output of the Linguistic Text Encoder, $F^{t}_{ling}$, acts as $Q$.
@@ 3.2 Multi-Scale Cross-Attention and Visual Decoder
Multi-Scale Cross-Attention. The collective use of Global and Local Cross-Attention ensures a balanced fusion of global emotional context and local linguistic features. As corroborated in Section 4, this multimodal fusion mechanism enables FT2TF to generate talking face videos with natural lip synchronization and emotionally expressive facial details, resulting in more realistic and engaging outputs.
Alignment. In the realm of text-video talking face synthesis, a primary concern revolves around the alignment between textual inputs and video frames. Given the variability in speech rates and temporal intervals between words, especially when considering text-only inputs, we address this challenge by circumventing the alignment problem between text and video frames. Instead, we leverage extensive audiovisual data alongside a highly efficient lip synthesis expert module to ensure the synthesis of text-video talking faces. Operating on an end-to-end basis, our model necessitates correspondence between the video input and the textual content, thereby facilitating the generation of lip images that closely approximate the ground truth during synthesis. As illustrated in Figure 3, the emotional content and latent linguistic information of the text also influence different facial regions through Multi-Scale Cross-Attention. This phenomenon indicates that our model has successfully aligned features from different distributions.
@@ 3.3 Pipeline Training
Three key loss functions are constructed for pipeline training.
Generation Loss. The generation loss, $L_{gen}$, measures the pixel-wise discrepancy between the generated frames and the ground truth ones, with the length represented by $N_{gen}$. This loss primarily focuses on supervising the accurate generation of pixel-level details to ensure the fidelity of the synthesized frames, as shown below:
$L_{gen}=\frac{1}{N_{gen}}\sum_{i=1}^{N_{gen}}\left\|X^{v}_{k+i}-O^{tv}_{k+i}\right\|_{1}.$ (1)
@@
$L_{syn}=\frac{1}{N_{gen}}\sum_{i=1}^{N_{gen}}SyncNet(O^{tv}_{k+i},A_{k+i}).$ (2)
Discriminator Loss. The discriminator loss [12], $L_{disc}$, aims to guide the encoder-decoder process to produce more realistic and coherent images. It operates by utilizing a discriminator that provides binary predictions $O_{pred}$ based on whether the input is the generated one or the ground truth. Specifically, the corresponding labels $Y_{disc}$ are assigned the label 1 for the generated data and 0 for the ground truth, producing a binary cross-entropy loss:
$\begin{split}L_{disc}=-\frac{1}{N_{gen}}\sum_{i=1}^{N_{gen}}(&Y_{disc}\cdot\log(O_{pred})\\ &+(1-Y_{disc})\cdot\log(1-O_{pred})).\end{split}$ (3)
@@
$L_{total}=\lambda_{1}L_{gen}+\lambda_{2}L_{syn}+\lambda_{3}L_{disc}.$ (4)
@@ 4 Experiment
4.1 Datasets
  • LRS2 [1]: The dataset comprises 48,164 video clips sourced from BBC outdoor broadcasts, divided into training, validation, and test sets with 45,839, 1,082, and 1,243 clips, respectively. Each clip is paired with a text caption of up to 100 characters. To ensure uniform input lengths, we retain videos with frame counts between 30 and 35.
  • LRS3 [2]: The dataset comprises over 400 hours of content from 5,594 indoor TED/TEDx talks, totaling 151,819 videos with corresponding text captions. Specifically, it includes 32,000 training/validation utterances and 1,452 test utterances. Consistent with the LRS2 dataset, we retain videos with frame counts between 30 and 35.
Although the two datasets share similar names, it is important to emphasize that they have no overlap. The LRS2 and LRS3 videos differ significantly in terms of scenes, data collection methods, character actions, environmental conditions, and other attributes.
4.2 Comparison Methods
We conduct a comprehensive evaluation of our method against state-of-the-art approaches for talking face generation. To demonstrate the effectiveness and efficiency of our model, we compare it with state-of-the-art audio-driven and text-driven models on the LRS2 and LRS3 datasets. These models are categorized based on two criteria:
  • Generation Task: Methods are categorized into either Lip Region or Whole Face, with the former generating only the lip region and the latter generating the entire face.
  • Modality: Methods are classified based on their input modality, including Audio+Video (A+V), Text+Video (T+V), Audio+Video+Pose/Landmarks (A+V+P/L).
4.3 Evaluation Metrics
We employ multiple evaluation metrics across various criteria, ranging from evaluating video synthesis quality to evaluating facial expression representation. To evaluate the similarity between generated results and ground truth ones, we apply Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) [41]. In addition, we utilize Learned Perceptual Image Patch Similarity (LPIPS) [47] and Frechet Inception Distance (FID) [14] for assessing feature-level similarity. Considering identity preservation, we leverage cosine similarity (CSIM) between identity vectors extracted by the ArcFace [8] network. Moreover, aligning with the approach outlined in [42], we measure the normalized lip landmarks distance (LipLMD) between generated and ground truth frames to evaluate lip synchronization quality. The scores for all methods are calculated over the entire face for a fair comparison.
[…]